In an experiment you have 5 moles of Fe, how many grams of Fe2O3 can be produced?


1
Expert's answer
2021-05-21T02:34:11-0400

As we can easily observe that "Fe_2O_3" contains 2 moles of "Fe" .


Hence 2 mole of "Fe" "= 1" mole of "Fe_2O_3"

1 mole of "Fe= \dfrac{1}{2}" mole of "Fe_2O_3"

5 moles of"Fe=" "\dfrac{5}{2}" moles of "Fe_2O_3"


Now:

Weight of "Fe_2O_3" = moles of "Fe_2O_3 \times" molar mass of "Fe_2O_3"


Weight of "Fe_2O_3=" "\dfrac{5}{2} \times 159.69=399.225 gm \approx 400 gm"